A simple 3 phase 2 port symmetrical π network has a series impedance of 10∠30⁰, and each shunt admittance 2∠30⁰ mS. The network has an output load of 50MW at 66kV with 0.9 lagging power factor.
Given that , calculate:
a) The ABCD parameters
b) The voltage, current and power factor at supply side
Answer #1

. RR VR 10/30 m 2 ANA * vos locais on / 4 1 2 = 2x103 w/30 2x10 SA <30 Y= 4x10°/30 t A= D= (1+ y) =(1+ 40x10*/60 ) = 1:0!014.Is = CVR TDIR = (4.02x1033049) 66 + (1-01014-0.9824) 0-48998 1-25-84) Is= 0.591 (-12-550 coop = coo ( 12.55 +1.338) = 0.97 la